# Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment

### v1

@@ 5.2 LLM-based Approaches
These approaches do not rely on trial-specific heuristics. Yet, as systems are evaluated on a few disease-specific trials, the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al. (2023), the generalizability of LLMs across diseases is unclear. Moreover, when compared, LLMs are less effective in making trial-level predictions (low F1 scores) than criterion-level ones.
@@
Wong et al. (2023) used GPT4 to structure eligibility criteria into logical expressions. Despite lagging behind expert systems, which use heavy manual intervention, they demonstrate the readiness of general-purpose LLMs for cohort-specific NER (oncology in this case) and logical structuring. Zero-shot and few-shot LLM prompts outperformed the state-of-the-art biomedical NER, SciSpaCy Neumann et al. (2019).
Yuan et al. (2023) prompt LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then composed by asserting that all inclusion criteria and none of the exclusion criteria match. On six trials for stoke patients, the model achieved 0.91 F1 at criterion-level vs. 0.81 F1 at trial-level.
Criterion-Level Prediction
Hamer et al. (2023) use OpenAI’s text-davince-003 model to first filter criteria that applies to a patient, then justify whether applicable criteria are met and to finally predict criterion-level labels based on previous explanations. This effectively reduces the total number of eligibility criteria (down to 300 from originally 4,100) a physician has to assess before recommending trials to a patient.
Wornow et al. (2025) prompt LLMs with trial-patient pairs to generate structured JSON strings of criterion-level predictions, explanations with evidence and a self-declared confidence. The content of the prompt ranges from the entire patient record and a trial to a single criterion and relevant record snippets. GPT4 achieves a micro-F1 of 0.93 vs. 0.91 points by the prior state-of-the-art on the 2018 N2C2 cohort dataset. The RECTIFIER Unlu et al. (2024) employs a RAG-based approach to find top-k relevant patient data stored as vector embeddings. The criteria and relevant records are prompted to the LLM to generate either Yes or No. Evaluated on just a single trial and a pool of 3000 private patient records, with clinical notes spanning three years, RECTIFIER achieved an accuracy higher than trained non-licensed staff overall (92.7% vs. 89.1%), especially for the Symptomatic Heart Failure criterion (97.9% vs. 91.7%).
@@
One line of work investigates the role of neural re-rankers for ranking relevant trials. Pradeep et al. (2022) generate multiple independent queries from a patient description to retrieve relevant trials. The trial and patient description is then passed to a fine-tuned T5 model to generate a relevance label. Zhuang et al. (2023) propose a multi-stage pipeline that involves a hybrid sparse-dense retriever, initialized with PubmedBERT-base and fine-tuned to retrieve the top 1000 trials. A cross-encoder, initialized with PubmedBERT-large, is trained to re-rank these results. Finally, the top-20 trials are re-ranked using GPT4. Rybinski et al. (2024) leverage LLMs for query reformulation and trial re-ranking to outperform traditional BM25 and BERT-based retrieval models. They show that BM25 is an effective initial retriever, surpassed by BERT-based neural re-rankers.
Jin et al. (2024) use GPT4 to first extract query keywords to retrieve top-k trials and then prompt GPT4 with patient data and eligibility criteria to generate criterion-level labels and explanations. They use a combination of manually-defined and GPT4 prompted scores to determine a trial-level score for final raning. Their TrialGPT performs better on the TREC datasets compared to smaller domain-specific transformer models and show a screening time reduction by 42%. Trial-LLama by Nievas et al. (2024) extend this study by extensive fine-tuning of open-source LLaMA models.
6 Discussion
Annotated Corpora
There have been few attempts to analyze and structure the clinical trials repository. As a result, finding similar trials and analyzing systematic failure cases is notoriously difficult. Chia, a manually annotated corpus of 1000 clinical trials created by Kury et al. (2020) has criterion-level annotations of entities, relations and Boolean logic graphs. Such annotations on trial-patient matching benchmarks can help us measure model fairness. Scaling patient data collections is also challenging as evident from Table 1, where patient records are in the hundreds. Individual records, curated manually, range from keywords via short patient summaries to documents over a time period. The format ranges from free text via semi-structured in-line tables to full-fledged databases. The rich and large MIMIC dataset, remains largely untapped for the task of trial-patient matching. LLMs provide a promising avenue for solving the automatic annotation at scale.
Grouth-truth Labels
Access to trial-patient eligibility labels, i.e., {eligible, ineligible, not-relevant}, is integral for training and evaluation. As they are obtained from human experts, they are difficult to scale and take time. Hence, many systems are evaluated retrospectively on patient data from private repositories such as clinics that maintain participation records in past trials Wong et al. (2023); Yuan et al. (2023); Hamer et al. (2023); Unlu et al. (2024). The largest public annotations from the TREC CT tracks have an average of 700 trial annotations per patient description. Bedi et al. (2025) found that of the 519 studies between 2022 and 2024 that evaluated LLM applications in healthcare, only 5% used real patient data. Jointly, the major public data sources, e.g., ClinicalTrials.gov and MIMIC datasets, present an exciting opportunity to develop generalizable matching methods.
Formalization
We formalize the task of trial-patient matching in Equation 3 through simple set operations. Surprisingly, except Yuan et al. (2023), none of the related work explicitly formalize the task. This coincides with a lack of aggregation strategies at the criterion-level Hamer et al. (2023); Wornow et al. (2025); Unlu et al. (2024). Formalizing the task, guides Yuan et al. (2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria. The criterion-level labels by Yuan et al. (2023), while not being directly interpretable, are aggregated as in Equation 3 for the trial-level labels. Wong et al. (2023) could potentially explain prediction via logical component mismatches, though the current F1 measures are below 50 points. Clinical experts, who analyze matches for the final cohort selection, strictly follow Equation 3. Criterion-level predictions and explanations have shown to reduce screening time by 42% Jin et al. (2024).
Core Limitations of LLMs
According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7, we discuss how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
Societal Impact
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al. (2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al. (2024). There is a lack of such analyses and mitigation methods in trial-patient matching and a lack of diverse datasets. The popularity of propreitary models for biomedical tasks Lu et al. (2024) including trial-patient matching Wornow et al. (2025); Jin et al. (2024) further ethical concerns on handling sensitive patient data. Alber et al. (2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. More research is required to determine the feasibility of locally deploying open-source models, the associated costs of maintaining them and keeping them up-to-date.
@@ 6 Discussion
7 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we come across semantically overlapping error types that are used to evaluate LLM generated explanations. For instance, incorrect reasoning and lack of knowledge are recognized as independent error types in Jin et al. (2024); Hamer et al. (2023); Nievas et al. (2024), even though the latter often leads to the former. In the proposal by Liévin et al. (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. We therefore identify where errors occur and break them down further by an identifiable source of the error (Figure 2).
@@ 7 Taxonomy of LLM Errors
Any error in the explanations provided by LLMs fall into this category. The most common source is insufficient explicit data. Basically, LLMs fail to draw logical conclusions from given data, while an expert can. This stems from previously unseen data (“lack of knowledge”) or the inability to recall prior information (“lack of implicit reasoning”) or the inability to infer from context (“insufficient data”).
The second source of reasoning errors is lack of self-assessment. The LLM contradicts explicit information in the prompt. The error occurs when knowledge is wrongly recalled or knowledge is correctly recalled, but contextual information is not applied, resulting in wrong reasoning. This is often referred to as “incorrect knowledge”.
The third source are premature decisions made by an LLM when the criteria actually require information that is missing, such as, expert opinion.
@@
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al. (2022). Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al. (2024)). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to report unless explanations for correct predictions are evaluated as well. It may seem as if the LLM is taking a shortcut, hinting towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
@@ 8 Future Directions
Clinical trial recruitment spans multiple interactive parts. As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
@@
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This task goes beyond keyword search to match eligibility criteria of two trials. This problem has seen little activity since clustering using lexical features Hao et al. (2014). Analogous systems, such as SemEHR Wu et al. (2018), exist for interactive semantic search over patient EHRs. Sun et al. (2022) combine multiple similarity measures to account for categorical features (study type, phase, and so on) and lexical features (condition name, intervention name, outcome measures, eligibility criteria) to find similar trials. Jullien et al. (2023) deploy LLMs to find trials that match short descriptions using textual entailment and ranked evidences. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
@@
Investigators design trials independent of patient data. This creates a semantic gap between eligibility criteria and patient data. LLM agents have the potential of bridging this gap by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new Luo et al. (2013), however, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al. (2017); Sun and Loparo (2019); Liu et al. (2021a); Dasgupta et al. (2020). Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022). LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al. (2021b).
@@
Eligibility criteria have a range of difficulty levels. A preemptive task can interact with patient records to estimate the size of the patient pool on the most restrictive criteria, say the presence of a rare genetic disorder. As patient records progressively follow uniform data models, it is more efficient to separate and structure criteria that require tabular operations (e.g., via structured queries) from criteria that require reasoning (temporal, numerical, negation), and to separate criteria that require expert analysis from being wrongly used in semantic searches Mozannar and Sontag (2020). Markey et al. (2025) showed promising results on content relevance and suitability with room for improvement in logical reasoning and provenance when using LLMs to generate trial protocols. Training LLM agents to identify and distribute tasks and then aggregating is a fertile ground for innovation.
@@
In the context of LLMs, accurately explaining black-box predictions in human-understandable form becomes very challenging Zhao et al. (2024a). Explainability research in clinical trial matches has been limited to prompting LLMs and manually evaluating chain-of-thought generations post-hoc, which is only one of the many facets of evaluating explainability Nauta et al. (2023); Bodria et al. (2023); Chen and Eickhoff (2024). Moreover, it is unclear how helpful these explanations are, especially, as discussed in Section 7, since they give rise to inconsistency errors and are categorized into ambiguous error labels.

### v2

@@ 6.2 LLM-based Approaches
Unlike classical approaches Penberthy et al. (2010); Zhang and Demner-Fushman (2017); Yuan et al. (2019); Tun et al. (2023), LLM-based approaches do not rely on trial-specific heuristics. Table 2 lists the primary differences between the classical and LLM-based approaches (see Appendix C for a full comparison). We group the LLM-based approaches by the granularity of the matches, starting with criterion-level prediction via trial-level prediction to trial ranking. While criterion- and trial-level predictions are direction-agnostic, trial ranking is a patient-centric task. With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in Ni et al. (2015)’s work, though this has not yet been addressed.
Criterion-Level Prediction
Here, methods utilize the reasoning capability of LLMs to obtain decision rationale and other context data in addition to the eligibility label. Hamer et al. (2023) use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria. Unlu et al. (2024); Beattie et al. (2024) and Wornow et al. (2025) chunk longitudinal patient data and store them as vector embeddings. They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with Beattie et al. (2024) providing expert criterion-level strategy in the prompts. Unlu et al. (2024) generate only the decision label from GPT4 and Beattie et al. (2024) and Wornow et al. (2025) generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context. Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and Wornow et al. (2025) additionally used open-sourced Llama Jiang et al. (2023) and Mixtral models Jiang et al. (2024). Ferber et al. (2024) prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
@@
Wong et al. (2023) used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information. Yuan et al. (2023) use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then computed by asserting Equation 6.1.
@@
Here, methods formulate effective queries, retrieve trials and re-rank them. Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model Pradeep et al. (2022) or zero-shot LLM prompts Saeidi et al. (2023); Kusa et al. (2023b), generating patient descriptions from trial data via 1-shot LLM prompts Zhuang et al. (2023), generating no-SQL queries via LLMs Ferber et al. (2024), using LLMs to reformulate and expand queries Rybinski et al. (2024); Peikos et al. (2024); Datta et al. (2025), and using LLMs to extract keywords Jullien et al. (2024); Jin et al. (2024); Nievas et al. (2024). This is followed by retrieval, either using embeddings similarity only Lahiri et al. (2023); Richmond and Deshpande (2023); Saeidi et al. (2023); Ferber et al. (2024); Saeidi (2025) or a multi-stage retrieval with neural re-rankers Zhuang et al. (2023); Rybinski and Karimi (2023); Rybinski et al. (2024); Jin et al. (2024); Datta et al. (2025). Certain approaches re-rank the top trials using GPT models Zhuang et al. (2023); Rybinski et al. (2024); Datta et al. (2025), some prompt LLMs for relevance labels Pradeep et al. (2022), while others prompt LLMs to generate trial- or criterion-level eligibility labels Rybinski et al. (2024); Peikos (2023); Jin et al. (2024); Nievas et al. (2024); Jullien et al. (2024). Criterion-level labels are aggregated using set-based reasoning mechanisms Jullien et al. (2024), variations of the matching Equation 6.1 Jin et al. (2024); Nievas et al. (2024); Saeidi (2025), and by prompting LLMs Jin et al. (2024).
7 Evaluation
Standard IR metrics, such as, precision, recall, F1, precision at rank k, normalized discounted cumulative gains (NDCG@k) and mean reciprocal rank (MRR), are the most popular evaluation metrics. The 2018 N2C2 cohort selection and TREC CT benchmarks have high performance metrics (Table 1), which are further improved by newer methods. Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data Kim and Quintana (2022); Kantor and Morzy (2024b).
Besides accuracy metrics, Rybinski et al. (2024) report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10. Jin et al. (2024) report a 42% reduction in screening time when using LLM-generated predictions and explanations. The measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent.
@@ 7 Evaluation
7.1 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we often come across inconsistent and semantically overlapping categories of errors in LLM generated explanations. For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by Jin et al. (2024); Hamer et al. (2023); Nievas et al. (2024), even though the latter often leads to the former (full list in Appendix B). In the proposal by Liévin et al. (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. Here, we identify where errors occur and break them down further by an identifiable source of the error (see Figure 2).
@@ 7.1 Taxonomy of LLM Errors
An error in the explanations provided by LLMs falls into this category. The most common source is insufficient explicit data, which occurs when LLMs fail to draw logical conclusions from given data, while an expert can. This stems from previously unseen data (“lack of knowledge”) or the inability to recall prior information (“lack of implicit reasoning”) or the inability to infer from context (“insufficient data”). The second source of reasoning errors is lack of self-assessment. The LLM contradicts explicit information in the prompt. The error occurs when knowledge is wrongly recalled or knowledge is correctly recalled, but contextual information is not applied, resulting in wrong reasoning. This is often referred to as “incorrect knowledge”. The third source are premature decisions made by an LLM when the criteria actually require information that is missing, such as, expert opinion.
@@
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al. (2022). Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al. (2024)). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to evaluate. This shortcut, similar to cognitive biases in humans, hints towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
8 Discussion
Annotated Corpora and their Size
Despite attempts to structure clinical trials Chen et al. (2022), finding similar trials and analyzing systematic failure cases is notoriously difficult Rybinski et al. (2020); White et al. (2023). Criterion-level annotations require significant manual annotations, thereby limiting the size of the corpus. For instance, Chia Kury et al. (2020) and LCT Dobbins et al. (2022), are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE Kang et al. (2017) and Criteria2Query Yuan et al. (2019), require expensive manual labels (230 disease-specific clinical trials in this case). Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy Kim and Quintana (2022). Systems thus evaluated on private retrospective data Wong et al. (2023); Yuan et al. (2023); Unlu et al. (2024) cannot be transparently compared.Kantor and Morzy (2024b) stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies. Public annotations, such as the TREC CT tracks have an average of 700 trial annotations per patient for less than 100 patients, while the N2C2 has only a few thousands of criterion-level annotations. Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by Kim and Quintana (2022) to generate large-scale data using automated methods.
Dimensions of Evaluation
The benchmarks in Table 1 focus on model accuracy, corroborating the result from Bedi et al. (2025), which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured. Omar et al. (2024) reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined. Further results from Nemati et al. (2025), who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
Formalization
Equation 6.1 highlights the importance of aggregation and priority. Surprisingly, very few works explicitly formalize the matching task. This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions Hamer et al. (2023); Wornow et al. (2025); Unlu et al. (2024) and others. Formalization guides Yuan et al. (2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, Jullien et al. (2024) and Saeidi (2025) to define re-ranking scores.
Societal Impact
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al. (2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al. (2024). Alber et al. (2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al. (2023), making the generalizability of LLMs across diseases unclear.
Data Sensitivity
While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations Unlu et al. (2024); Wong et al. (2023); Wornow et al. (2025), processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance Edemekong et al. (2024). As already discussed, creating large-scale realistic patient records while protecting their privacy is particularly challenging. There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative Das et al. (2023); Wang et al. (2024).
Core Limitations of LLMs
According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7.1, we discussed how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
@@ 9 Future Directions
As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
@@
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This problem has seen little activity since clustering using lexical features Hao et al. (2014). Newer search methods include constructing a clinical trial knowledge graph Chen et al. (2022), searching via patient EHRs Wu et al. (2018) and designing features for similarity matchingSun et al. (2022). Jullien et al. (2023) use textual entailment in LLMs to find trials that match short descriptions. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
@@
LLM agents have the potential of bridging the semantic gap between eligibility criteria and patient data by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new Luo et al. (2013), yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al. (2017); Sun and Loparo (2019); Liu et al. (2021a); Dasgupta et al. (2020). Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022). LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al. (2021b).
@@
Liu et al. (2025c) propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction. Markey et al. (2025) showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance. Similar to Shi et al. (2024), who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result. Another important operation is learning to defer to experts Mozannar and Sontag (2020), which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
@@
Explaining black-box LLM predictions in human-understandable form is very challenging Zhao et al. (2024a). Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability Nauta et al. (2023); Bodria et al. (2023); Chen and Eickhoff (2024). Wong et al. (2023) could potentially explain eligibility prediction via logical component (mis)matches. Furthermore, the utility of these explanations are unclear, as they have been only evaluated manually. We hope that the error taxonomy discussed in Section 7.1 assists in systematic evaluation of explanations.